UT Sep 10, 2019 - Restart Night #9


OSA: S. Allanson

ISp: C. Bolyard

Mtn Manager: M. Wagner

ISA: B. Rothberg, O. Kuhn (Tucson)

AO; D. Miller (Tucson)

Software: P. Kubanek

Summary:
  • LUCI Flexure collection
    • G210-N180 verification
    • N30-imaging collection
  • MODS testing
    • MODS GUI offsets
    • Calibrations

Plan Outline:

Plan rolled over from Night #7-8
  • LUCI
    • AGW-1 Transform
      • No need to update PCS, rotator ZP looks good after astrometric mask test
    • IF spectroscopic modes are released for LUCIs
      • G200 science-like OB
      • G210 science-like OB
  • Test patch to GCS/PCS to address dithering NS guiding
    • LUCI
    • MODS
  • DX M3 alignment
  • LBC
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters
    • LBC sub-framing (test OBs needed)
    • NS guiding with dithers
    • Checkout observations with Bessel filters
  • MODS
    • Test NS guiding with dithering
    • Test of MOD SGUI update coordinate type and offsets


-C. Bolyard cleaned g-sloan filter

UT 01:25 - LUCIs up and test exposure taken, troubleshooting earlier on hung LBC processes - seems resolved. PCS troubleshooting being done with S. Allanson and P. Kubanek

UT 01:36 - raining on mountain

UT 01:55 - S. Allanson checked outside - definitely wet, sense of impending doom - but not ready to call just yet

UT 02:49 - building still soaked, humidity climbing, nice thick clouds rolled overhead

UT 03:57 - O. Kuhn took a series of LBC biases, J. Hill testing some M3 motions

UT 04:22 - hit 90% humidity

UT 04:58 - authorizing MODS (although LUCI's in the beams) - O.Kuhn testing MODS offset issue in the MODS GUI
  • Sending a track mode preset which will make a detxy absolute offset immediately after the preset and then pause.
    • At pause, send detxy rel offset from gui (offsetxy -4 2.3 rel). ISISlog reports: offsetpointing 0 -4 2.3 DETXY ABS left
    • At pause, on right, toggle the detxy and rel buttons before sending offset (offsetxy -5 -3.4 rel). ISISlog reports: offsetpointing 0 -5 3.4 DETXY REL right
    • So this confirms that you must toggle the buttons on the gui. No need to switch from REL to ABS and back to REL, just click on REL.
  • Sending a track mode preset which makes an radec rel offset immediately after the preset.
    • Now the GUI reports that the offset coordsys is RADec.
    • Sending a rel radec offset of -3, -5.4 on MODS1: RADEC Rel (according to GUI). ISISlog reports: offsetpointing 0 -3 -5.4 RADEC REL left
    • Sending an abs radec offset of 2, 3.2 on MODS2: RADEC Abs (Abs selected on GUI). offsetpointing 0 2 3.2 RADEC ABS right
  • Sending a track mode preset which makes an absolute RADEC offset. Before sending, MODS1 GUI showed RADEC REl and MODS2 GUI showed RADEC Abs (because I had toggled the button).
    • Sending a REL RADEC offset to MODS1: offset 2.4 -1.2 rel: ISISlog reports offsetpointing 0 2.4 -1.2 RADEC ABS left
    • Sending an ABS RADEC offset to MODS2: offset 3.5 -2.3 abs: ISISlog reports offsetpointing 0 3.5 -2.3 RADEC ABS right
  • Sending a track mode preset with a rel radec offset. Will the button on the MODS2 GUI that was toggled to Abs change back to Rel? No, it remains as Abs.
  • One more time - this time sending a detxy rel. This caused the CoordSys on the GUI to change from RADec to DETXY, but not the MoveType.
  • Conclusion is that the CoordSys button on the GUI updates according to the type of offset being sent by the script, but the MoveType button does not.


UT 05:47 - flexure verifiation tests for N180 G210 - Argon cal lamp, J-band filter, Optic Sieve mask at 79 and 35 elevations

-configuring LUCI-1 MOS to optic sieve

UT 05:49 -LUCI-2 MOS error immediately
  • storage to FPU - Optic Sieve Slot 0
    • storage to turnout
    • MHU translator to mask
    • Not True that MHU is at required position
    • MHU angle should be 185.4, current is 186.06 - storage angle tolerance 1.2degrees
    • MHU position should be 0, current is -12
    • Elevation 79, rotator 210
    • Rotator is at 341 - ok
    • failed post condition check so not where its supposed to be MHU translator
    • grabber neg limit active so nothing grabbed, retainer shows all maske dlocked (card2 motor2 positive on)
    • initialize MHU translator is recover - nothing is grabbed
    • proceeding to initialize MHU
      • spindle is now at -128, angle 310
  • 2nd MOS error - missed grab on mask
    • MHU picker arms failed to grab mask
    • masks still locked
    • grabber moved 8000 steps, then 500, then 1000, straing gauges are on, but the limit switch is not
    • Recovery:
      • 1.1 MOS MCU set 0 counter
      • move -8000 to hit negative limit switch on grabber
      • negative limit switch came on, it moved -5118 steps
      • check preconditions to initialize - grabber - pre-conditions are OK - init
      • this was the wrong set I should have hit Translator initialize - DO NOT MAKE THIS MISTAKE
      • NOW initiailzing the MOS UNIT (this was the correct step)
      • Turning HC off on rotator
  • Recovered MOS errors
UT 06:20 - putting in cal unit - we hit 95% humidity

UT 06:23 - Argon w/J - flex comp is on

UT 06:24 - starting Elev 79

L2 0072-0073 IGNORE THESE

UT 06:26 - got a left rotator failed on SX

-probably timed out - need to manually move to -90 for both

-manually moving to -90 on DX to syn things up - G210 CenWave = 1.25

UT 06:28 - NOW restarting
  • EL 79
    • L1 0074-0094
    • L2 0074-0094
UT 06:51 - moving to El 35, staring at UT 06:52
  • EL 35
    • L1 0095-0115
    • L2 0095-0115
UT 07:15 - verifiation run finished

Reconfigure for N30+MIRR Calib15 with Halo1 - flexure off, J-band, OptSpec sieve - Elevation: 90,68,46,24

-already at rotaors=-90 from end of last script

L1/L2 0116 - test exposure - counts look OK

UT 07:23 - starting
  • El 90
    • L1 0117-0189
    • L2 0117-0189
  • EL 68 - UT 07:59
    • L1 0190-0262
    • L2 0190-0262
  • EL 46 - UT 08:38
    • L1 0263-0335
    • L2 0263-0335
  • EL 24 - UT 09:14
    • L1 0336-0408
    • L2 0336-0408
UT 09:50 - finished flex cals

-moving back to EL 90 and park positions for LUCI-1 and LUCI-2

put blind filters in, turned off cal and parked it - taking dark exposures to subtract for data

L1 0409-0410 - 5sec dark frame

L2 0409-0410 - 5sec dark frame


MODS Calibrations:

Running scripts from MODS Library after editing titles of some to reflect actual config.
  • grpixflat_r
    • execBinoMODS MODS1fix/grpixflat_r.txt1 MODS2fix/grpixflat_r.txt2
    • M1R: 2-6 and M2R: 2-6
  • grpixflat_b - first round for MODS1 had too-high counts and I dropped exptime from 3->2 for both Clear & UG5.
    • M1B: 2-6 (clear) and 7-11 (UG5) and M2B: 2-6 (clear) and 7-11 (UG5)
  • grpixflat_b - repeat with edited script for MODS1
    • execBinoMODS MOSD1fix/grpixflat_b.txt1_edit MODS2fix/grpixflat_b.txt2
    • M1B 12-16,17-21 & M2B 12-16,17-21
  • prlamps
    • execBinoMODS MODS1fix/prlamps.txt1 MODS2fix/prlamps.txt2
    • M1B 22-25 and M1R 7-10
    • M2B 22-25 and M1R 7-10
  • prlamps_b
    • execBinoMODS MODS1fix/prlamps_b.txt1_edit2 MODS2fix/prlamps_b.txt2_edit2
    • edited exptimes for Kr and Xe by factor 10 to avoid super-saturation wrap-around of lines at red end where resolution is lower
    • M1B 26-29 and M2B 26-29
  • prlamps_r
    • execBinoMODS MODS1fix/prlamps_r.txt1 MODS2fix/prlamps_r.txt2
    • M1R 11-14 and M2R 11-14
  • prslit1.0:
    • execBinoMODS MODS1fix/prslit1.0.txt1 MODS2fix/prslit1.0.txt2
    • M1B 30-34,35-39 and M2B 30-34,35-39. The ND1.5 are OK, but the UG5 (35-39) are mildly sat, the best that could be done.
    • Uses qth, ND1.5 and 0.5 sec.
  • prslit1.0_b:
    • execBinoMODS MODS1fix/prslit1.0_b.txt1 MODS2fix/prslit1.0_b.txt2
    • M1B 40-44, M2B
    • MODS2 script uses QTH instead of VFLAT. M2B 40-45 sat and M2B 46-47 super-sat. Control-c it and run the MODS1 script on MODS2
      • M2B 48-52 (30k) and 53- (23k)
  • bin12 grpixflats:
    • execBinoMODS MODS1fix/grpixflat_bin12.txt1_edit4
    • M1B 50-54 (22k) and 55-59 (UG5, 30k) and M1R 20-24 (20k)
    • M2B 58-62 (15k) and 63-67 (UG5, 24k) and M2R 20-24 (25k)
    • apparently vflat is weaker on mods2. Could increase time or lamp level for mods2.
  • bin12 grpixflats with custom mods2 script to increase Clear exptime from 4 to 7.
    • M1B 60-64 (27k) and 65-69 and M1R 25-29 (23k)
    • M2B 68-72 (32k) and 73-77 and M2R 25-29 (25k)

LBCs are turned off (really... no hung UI like yesterday) and MODS1+MODS2 are in bed.

-- BarryRothberg - 10 Sep 2019
Topic revision: r7 - 10 Sep 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback